OpenStudy (anonymous):

find the distance between (4/5, -1/2) and (2,0). express answer in simplest form.

OpenStudy (anonymous):

\[13/ 10\]

OpenStudy (angela210793):

A (4/5, -1/2) and B(2,0) AB=\[\sqrt{(Xb-Xa)^2+(Yb-Ya)^2} \]

OpenStudy (angela210793):

All u have to do is replace Xb = 2 Xa=4/5 and Yb=0 Ya=-1/2